In how many years will it take for money at least quadruple at 12% compounded semi-annually?
Round UP your answer for n to the next whole number then convert to years to one decimal place.

Answer by lwsshak3(11628)

In how many years will it take for money at least quadruple at 12% compounded semi-annually?
Round UP your answer for n to the next whole number then convert to years to one decimal place.
**
Compound Interest Formula: A=P(1+i)^n, P=initial investment, i=interest rate per period, n=number of periods, A=amt after n periods.
For given problem:
A/P=4
i=.12/2=.06
n=number of compound periods
..
4=(1+.06)^n
4=(1.06)^n
log(4)=nlog(1.06)
n=log(4)/log(1.06)
n≈23.79 compound periods
or
n≈23.79/2≈11.9 yrs